I have two question one is what is 2^square root of 5y ^3
the second is what is x+y=12 if x= 2y what is x

and if you could give me examples or a website to help me study with

Answer by Alan3354(69443)

what is 2^square root of 5y ^3
----
Not clear.
===========================
the second is what is x+y=12 if x= 2y what is x
---
Poor wording, using "what is"
---
x+y=12
x= 2y
Solve for x
===================
x+y=12
x= 2y
2y + y = 12
3y = 12
y = 4
-------
x = 2y
x = 8